The NetBSD Project

Chris Demetriou

January 17, 1994

The NetBSD Project

Chris Demetriou

January 17, 1994

What is NetBSD?

• A free and complete BSD operating system

• Derived from Berkeley’s Net/2 release

• Runs on multiple hardware platforms

• “Supported” by volunteers

NetBSD’s Goals

• Provide a free, complete, stable BSD system

• Improvements, where necessary

• “Support” for popular platforms

• Avoid encumbering licenses Summary:

Provide a free system that professionals and hobbyists can use for whatever they want.

SLIDE 2 XXX (notes)

2-1

User-land Improvements

• Shared libraries!

• Approaching POSIX.2

• Filled in missing commands

• Keeping software up-to-date We’re focused on making NetBSD a “solid” system.

SLIDE 3 XXX (notes)

3-1

Kernel Improvements

• More architectures supported and enhanced

• Many file system-related improvements

• Approaching POSIX.1

• Networking improvements

• “Broken” kernel subsystems replaced or fixed SLIDE 4 XXX (notes)

4-1 The NetBSD Project, January 17, 1994 5

“Supported” Platforms:

• i386/i486

• Amiga

• HP 9000/300

• Macintosh (with 68020 and MMU or 68030)

• Sparc (sun4c) Some of the ports in progress:

• DECstation

• Sun3 SLIDE 5 XXX (notes)

5-1

Additions and Improvements to the File Systems

• Many bug fixes to UFS

• Split NFS server and client code

• ISO-9660 file system (with Rock Ridge extensions)

• MS-DOS file system

• Union mounts

• /proc, and other “miscellaneous” file systems SLIDE 6 XXX (notes)

6-1

Networking improvements:

• Multicast

• PPP Work done on kernel subsystems:

• New buffer cache, execve(), physio()

• Lots of tty-handling fixes

• Ring buffers with CLIST interface SLIDE 7 XXX (notes)

7-1

Compatibility with other systems

• BSDI binary compatibility

• SunOS binary compatibility

• POSIX

• Nearly 4.4BSD source-code compatible It’s easy to port software to NetBSD.

SLIDE 8 XXX (notes)

8-1

Future work on NetBSD

• POSIX conformance

• Further system integration

• 4.4BSD “Lite” – We’re still waiting!

• Better SCSI and autoconfiguration support

• Remote debugging support

• More platforms

• More file system work SLIDE 9 XXX (notes)

9-1

Who does the work?

A “cast of thousands” of contributors including: • CSRG, for Net/2 and for “BSD” in general

• The NetBSD core team

• Users who have sent in bug reports and fixes

• The “NetBSD-current” alpha-testers We’re not going away any time soon.

SLIDE 10 XXX (notes)

10-1

How to get NetBSD via anonymous FTP:

    agate.berkeley.edu [128.32.155.1]
    pub/NetBSD

NetBSD mailing lists:

    majordomo@sun-lamp.cs.berkeley.edu

And if you have questions or comments:

  • netbsd-comments@sun-lamp.cs.berkeley.edu

• Come talk to us privately

• Ask questions now!

SLIDE 11 Also: • NetBSD-current: updated nightly • SUP of current sources possible;

  sun-lamp:pub/sup for details

11-1